OpenStudy (anonymous):

Mia is three years older than twice her sister Brooke's age. The sum of their ages is less than 30. What is the greatest age Brooke could be? (Points : 1) a.)7 b.)8 c.) 9 d.)10

OpenStudy (anonymous):

I think 8

OpenStudy (anonymous):

y

OpenStudy (anonymous):

Let Mia's age be m and her sister's age be x m=2x+3 2x+3+x=less than 30 For greatest age we need to take largest number. 3x+3=29 x=26/3 But the options are all in whole numbers, so if 3x+3=27 then x =8

OpenStudy (anonymous):

LET HER SIS AGE BE x then mia=3+2x....x+(3+2x)=30 therefore....3x+3=30..... 3x=27...x=27/3.... x=9

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

@ Trustted Sum of their ages is less than 30

OpenStudy (anonymous):

i didnt c that... so u myt b correct